# coverforest: Conformal Predictions with Random Forest in Python

### v1

@@ Abstract
We present coverforest, a Python package that implements efficient conformal prediction methods specifically optimized for random forests. coverforest  supports both regression and classification tasks through various conformal prediction methods, including split conformal, CV+, Jackknife+-after-bootstrap, and adaptive prediction sets. Our package leverages parallel computing and Cython optimizations to speed up out-of-bag calculations. Our experiments demonstrate that coverforest’s predictions achieve the desired level of coverage. In addition, its training and prediction times can be faster than an existing implementation by 2–9 times. The source code for the coverforest is hosted on GitHub at https://github.com/donlapark/coverforest.
@@ 1 Introduction
Classification is a fundamental machine learning task with applications ranging from medical diagnosis to customer segmentation and computer vision. While numerous classification algorithms exist in the literature, most implementations focus solely on point predictions. However, in many real-world applications, particularly those involving critical decision-making, understanding the uncertainty associated with these predictions becomes crucial. For this reason, conformal prediction techniques have recently garnered significant interest, as they leverage data exchangeability to produce prediction sets/intervals that contain the true output with a guaranteed probability above a specified level.
Recent conformal prediction methods like CV+ and Jackknife+-after-bootstrap (J+ab) rely on out-of-fold and out-of-bag calculations—techniques extensively studied in ensemble learning, particularly in random forests [1]. Years of development in open-source machine learning packages have led to highly optimized out-of-bag calculations in random forests. The scikit-learn’s implementation, for instance, leverages parallel computing to speed up the calculations. Even though these implementations are only for in-sample out-of-bag calculations, we can adapt the technique for train-test cross-conformity score calculations, enabling faster conformal predictions with CV+ and J+ab.
Building on this idea, we introduce coverforest, a Python package for conformal predictions using random forest. coverforest implements several recent advances in conformal prediction, including split conformal, CV+ and J+ab for regression, and adaptive prediction set (APS) with split conformal, CV+, and J+ab for classification. Random forest is our model of choice for three reasons 1) they have consistently demonstrated good performance on tabular data [2], 2) they can be trained with missing data, and 3) trees are very fast to train in parallel, making them attractive for cross-conformal predictions like CV+ and J+ab.
In our implementation, trees are fitted and queried for predictions in parallel using the joblib [3] package in Python, and the cross-conformity scores are calculated in Cython [4]. Experimental results show that coverforest’s conformal training and prediction of random forests are faster than an existing implementation by 2–9 times faster.
@@ 2 Background on Conformal Predictions
Let $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ be training data for a predictive model $\hat{f}$ and $(X_{n+1},Y_{n+1})$ is a new test point, where $X_{i}$’s lie in an input space $\mathcal{X}$ and $Y_{i}$’s lie in an output space $\mathcal{Y}$. Without loss of generality, we assume that $\mathcal{Y}=\mathbb{R}$ for a regression task and $\mathcal{Y}=\{1,\ldots,C\}$ for a classification task. A prediction set for $Y_{n+1}$ is a set of outputs $\mathcal{C}=\mathcal{C}(X_{i})\subseteq\mathcal{Y}$ that might be input-dependent. For regression tasks prediction sets are usually restricted to intervals, in which case they are commonly referred to as prediction intervals.
In a conformal prediction task, we specify a miscoverage rate $\alpha\in(0,1)$, with which we aim to construct a prediction set $\mathcal{C}$ that has probability at least $1-\alpha$ to contain the true class:
@@
Classical results in conformal prediction literature typically assume that the random variables $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and $(X_{n+1},Y_{n+1})$ are either i.i.d. or exchangable, that is, their joint distribution are invariant under permutations. Below, we summarize relevant conformal prediction methods that provide prediction sets with coverage guarantees under one of these assumptions.
@@
    where $q^{-}_{\alpha}$ and $q^{+}_{\alpha}$ are defined as in CV+. However, this interval with a fixed number of bootstrap samplings $B$ does not achieve a finite-sample coverage guarantee, due to the asymmetry between training and test samples; to see this, observe that each test sample is excluded from exactly $B$ bootstraps, while each training sample is typically excluded from $<B$ bootstraps. This issue is resolved by making $B$ a random variable—by resampling $B$ from a Binomial distribution:
@@
    Adaptive Prediction Sets (APS) [8]. This method proposes the cumulative probability as the conformity score: given the predicted probabilities of a point $(x,y)$ in descending order $\hat{\pi}_{1}(x)\geq\hat{\pi}_{2}(x)\geq\ldots\geq\hat{\pi}_{C}(x)$, the score function is the (randomized) generalized inverse quantile:
    $E(x,y)=\hat{\pi}_{1}(x)+\ldots+\hat{\pi}_{y-1}(x)+u\cdot\hat{\pi}_{y}(x),\quad u\sim\operatorname{Uniform}(0,1).$ (4)
    In other words, the score is chosen randomly between the cumulative probabilities of $y$ and $y-1$. Without this randomization, the prediction sets would generally cover more than the desired level of ($1-\alpha$)if $E(x,y)$ is defined up to $\hat{\pi}_{y}$, and less if it is defined up to $\hat{\pi}_{y-1}$.
    For the split conformal, we let $q_{\alpha}$ be the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of the calibration scores. The prediction set for a new point $X_{n+1}$ is then:
    $\mathcal{C}(X_{n+1})=\left\{1,\ldots,k:k=\inf_{y\in\{1,\ldots,C\}}E(X_{n+1},y)\geq q_{\alpha}\right\}.$
    There is also a CV+ variant of this method: First, we split the data into $K$ folds. For each fold $\mathcal{I}_{k}$, we fit a model $\hat{\pi}^{(-k)}$ on all data except fold $k$ and then calculate the score $E(X_{i},Y_{i},\hat{\pi}^{(-S(i))})$ where $S(i)=k$ if $i\in\mathcal{I}_{k}$. The prediction set for a new point $X_{n+1}$ is then all $y$ such that less that $1-\alpha$ fraction of the training set has their scores less than the score of $y$ under the same model, in other words, $E(X_{i},Y_{i},\hat{\pi}^{(-S(i))})<E(X_{n+1},y,\hat{\pi}^{(-S(i))})$.
@@
    Regularized Adaptive Prediction Sets (RAPS). [9]. This method was proposed for classification with a lot of classes (say 1000+ classes) where the ordering of the predicted probabilities does not reflect the true order. To this end, the authors propose to add to the APS’s score function a regularization term that penalizes classes with small predicted probabilities (assuming $\hat{\pi}_{1}(x)\geq\hat{\pi}_{2}(x)\geq\ldots\geq\hat{\pi}_{C}(x)$ as before):
@@
    As with APS, we can use this score function for split conformal or CV+.
  3. 3.
    Jackknife+-after-Bootstrap (J+ab). This method extends Jackknife+after-bootstrap from regression to classification by using APS or RAPS score functions. Since the proof of the coverage lower bound applies to any exchangeable conformity scores, the lower bound remains valid when using these classification scores.
@@
APS + Split $1-\alpha$ $\approx 1-\alpha$ $O(Tdn\log n)$ $O(Tn_{\operatorname{test}}\log n)$
APS + CV+ $1-2\alpha-\epsilon_{K,n}$ $\gtrsim 1-\alpha$ $O(KTdn\log n)$ $O(Cn*n_{\operatorname{test}})$
APS + J+ab $1-2\alpha$ $\gtrsim 1-\alpha$ $O(Tdn\log n)$ $O(Cn*n_{\operatorname{test}})$
Table 1: Comparison of coverage guarantees and computational costs for different conformal random forest methods. Here, the notation $\gtrsim 1-\alpha$ means “sometimes worse than $1-\alpha$”, $n$ is the number of training samples, $n_{\operatorname{test}}$ is the number of test samples, $T$ is the number of trees, $C$ is the number of classes, $K$ is the number of folds in CV+, and $\epsilon_{K,n}=\min\left\{2(1-1/K)/(n/K+1),(1-K/n)/(K+1)\right\}$. We assume that $T\ll n$.
In Table 1, we compare coverage guarantees and computational costs across different conformal prediction methods for random forests. While CV+ and J+ab theoretically guarantee coverages of $\approx 1-2\alpha$, they typically achieve $\geq 1-\alpha$ coverage in practice. The methods with the APS score, however, have notably long prediction times due to the pairwise comparisons between train and test conformity scores across all classes.
@@ 2.1 Related work
The coverforest package implement conformal predictions [10, 5, 11, 12, 13] using random forest [1]. For similar work, conformity scores of bootstrapped regression based on iterative reweighted least squares is briefly discussed in [5]. For random forest classification, three conformity scores were proposed in [14], though their coverage guarantee held only with infinite data.
@@
For classification, given a model that outputs probability estimates $\hat{\pi}_{j}(x)\approx P(Y=j|X=x)$, a basic conformal prediction method can be devised based on a score function $s(x,y)=1-\hat{\pi}_{y}(x)$. The prediction set for a new point $X_{n+1}$ is then given by $\mathcal{C}(X_{n+1})=\left\{y:s(X_{n+1},y)\leq q_{\alpha}\right\}$, where $q_{\alpha}$ is the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of the calibration scores [22, 5]. However, this method tend to outputs an empty set when the model is uncertain about the prediction, for example, $q_{\alpha}=0.2$ and the predicted probabilities of two classes are 0.5 each. This issue can be remedied by instead using the cumulative probability as in the APS method [8]. Our future plans also include implementing multilabel classification with finite-sample coverage guarantee via conformal risk control [23, 24, 25].
@@ 3.1 API description
For moderately large training and test sets, calculating the conformity scores across all pairs of training and test samples can be quite expensive. To improve performance, the calculations have been optimized in Cython with parallel processing enabled. Users can specify the number of threads through the num_threads parameter in the predict() method.

### v2

@@ Abstract
We present coverforest, a Python package that implements efficient conformal prediction methods specifically optimized for random forests. Our package leverages parallel computing and Cython optimizations to accelerate out-of-bag calculations, making it 2–10 times faster than existing implementations while maintaining theoretical coverage guarantees. coverforest supports both regression and classification tasks through various conformal prediction methods, including split conformal, CV+, and Jackknife+-after-bootstrap, as well as adaptive prediction sets for classification. Built upon scikit-learn’s random forest implementation, coverforest provides a practical solution for reliable uncertainty quantification. We demonstrate the package’s performance through empirical evaluations on several benchmark datasets. The source code for the coverforest is hosted on GitHub at https://github.com/donlap/coverforest.
@@ 1 Introduction
Classification and regression are fundamental machine learning tasks with applications ranging from medical diagnosis to customer segmentation and computer vision. While numerous classification algorithms exist in the literature, most implementations focus solely on point predictions. However, in many real-world applications, particularly those involving critical decision-making, understanding the uncertainty associated with these predictions becomes crucial. For this reason, conformal prediction techniques [1, 2, 3, 4] have recently garnered significant interest, as they leverage data exchangeability to produce prediction sets/intervals that contain the true output with a guaranteed probability above a specified level.
Recent conformal prediction methods like CV+ [5] and J+ab [6] rely on out-of-fold and out-of-bag calculations, techniques frequently used in ensemble learning and particularly in random forests [7]. Years of development in open-source machine learning packages have led to highly optimized out-of-bag calculations in random forests. The scikit-learn’s implementation, for instance, leverages parallel computing to speed up the calculations. Even though these implementations are only for in-sample out-of-bag calculations, we can adapt the technique for train-test cross-conformity score calculations, enabling faster conformal predictions with CV+ and J+ab.
Building on this idea, we introduce coverforest, a Python package for conformal predictions using random forest. coverforest implements several recent advances in conformal prediction, including split conformal [4], CV+ [8] and J+ab [6] for regression, and adaptive prediction set (APS) [5] with split conformal, CV+, and J+ab for classification. Random forest is our model of choice for three reasons 1) they have consistently demonstrated good performance on tabular data [9], 2) they can be trained with missing data [7, 10], and 3) trees are very fast to train in parallel [11], making them attractive for cross-conformal predictions like CV+ and J+ab.
In our implementation, trees are fitted and queried for predictions in parallel using the joblib [12] package in Python, and the cross-conformity scores are calculated in Cython [13].
The following is a summary of coverforest’s contribution:
  1. 1.
    coverforest directly extends scikit-learn’s random forest, which means it also utilizes already well-optimized C+ code and parallel computation. Experimental results show that coverforest’s conformal training and prediction of random forests are 2–10 times faster than an existing implementation.
  2. 2.
    coverforest is designed to follow scikit-learn’s API conventions with the simple fit() and predict() methods. It can also be easily combined with scikit-learn’s preprocessing methods, allowing it to handle various data formats.
  3. 3.
    Even though there is an existing package that implements adaptive prediction sets (APS) for conformal prediction, coverforest is the first package to implement the Jackknife+-after-Bootstrap (J+ab) method for conformal prediction.
@@ 2 Background on Conformal Predictions
Let $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ be training data for a predictive model $\hat{f}$ and $(X_{n+1},Y_{n+1})$ is a new test point, where $X_{i}$’s lie in an input space $\mathcal{X}$ and $Y_{i}$’s lie in an output space $\mathcal{Y}$. Specifically, we let $\mathcal{Y}=\mathbb{R}$ for a regression task and $\mathcal{Y}=\{1,\ldots,C\}$ for a classification task. A prediction set for $Y_{n+1}$ is a set of outputs $\mathcal{C}=\mathcal{C}(X_{i})\subseteq\mathcal{Y}$ that might be input-dependent. For regression tasks we restrict prediction sets to intervals, in which case they are commonly referred to as prediction intervals.
In a conformal prediction task, we specify a miscoverage rate $\alpha\in(0,1)$, with which we aim to construct a prediction set $\mathcal{C}$ that has probability at least $1-\alpha$ to contain the true label:
@@
Classical results in conformal prediction literature typically assume that the random variables $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and $(X_{n+1},Y_{n+1})$ are either i.i.d. or exchangable, that is, their joint distribution are invariant under permutations. Below, we provide a review of well-known conformal prediction methods, and recently proposed conformity that achieve coverage guarantees under one of these assumptions.
@@
    where $q^{-}_{\alpha}$ and $q^{+}_{\alpha}$ are defined as in CV+. However, this interval with a fixed number of bootstrap samplings $B$ does not achieve a finite-sample coverage guarantee, due to the asymmetry between training and test samples; to see this, observe that each test sample is excluded from exactly $B$ bootstraps, while each training sample is typically excluded from <B<B bootstraps. This issue is resolved by resampling $B$ from a Binomial distribution:
@@
    Split Conformal with APS Score [5]. We split the data into training and calibration sets, then fit a classification model $\hat{f}$ on the training set. An obvious conformity scores are the model’s probability predictions; however, a hypothetical model that assigns 505050-50 probabilities to the most likely and second most likely classes would result in set predictions that only contain the most likely classes, and hence fails to achieve the target coverage level. Alternatively, Adaptive Prediction Sets (APS) [5] has been recently proposed method that is adaptive to data distributions.
    The crux of the APS method is the use of cumulative probability as the conformity score: given the predicted probabilities of a point $(x,y)$ in descending order $\hat{\pi}_{1}(x)\geq\hat{\pi}_{2}(x)\geq\ldots\geq\hat{\pi}_{C}(x)$, the APS score function is the (randomized) generalized inverse quantile:
    $E(x,y,\hat{\pi})=\hat{\pi}_{1}(x)+\ldots+\hat{\pi}_{y-1}(x)+u\cdot\hat{\pi}_{y}(x),\quad u\sim\operatorname{Uniform}(0,1).$ (4)
[…]
    In the case of split conformal prediction, we let $q_{\alpha}$ be the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of the calibration scores. The prediction set for a new point $X_{n+1}$ is then:
    $\mathcal{C}(X_{n+1})=\left\{1,\ldots,k:k=\inf_{y\in\{1,\ldots,C\}}E(X_{n+1},y,\hat{\pi})\geq q_{\alpha}\right\}.$
    The APS score typically produces larger prediction sets than the probability prediction score; however, the former generally achieves the target coverage level while the latter often fails to do so. In Section 5, we experimentally compare these two scores to confirm this observation.
@@
    Cross-Validation+ (CV+) with APS Score [5]. The CV+ method for classification aims to improve data efficiency compared to a simple train-calibration split by using a cross-validation scheme. This allows each data point to be used for both training and calibration, but never in the same role simultaneously. The procedure using the Adaptive Prediction Set (APS) score is as follows:
    1. (a)
      Data Partitioning. Split the training data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ into $K$ disjoint folds, denoted by $\mathcal{I}_{1},\ldots,\mathcal{I}_{K}$.
    2. (b)
      Cross-Validation Training. For each fold $k\in\{1,\ldots,K\}$, train a classification model $\hat{\pi}^{(-k)}$ on all data points except those in fold $\mathcal{I}_{k}$. This results in $K$ different models.
    3. (c)
      Leave-one-fold-out Score Calculation. We associate each data point $(X_{i},Y_{i})$, with the model $\hat{\pi}^{(-S(i))}$ where $S(i)=k$ if $i\in\mathcal{I}_{k}$ (i.e., the model that was not trained on the fold containing $i$) to calculate its APS conformity score. Given the predicted probabilities for $(X_{i},Y_{i})$ sorted in descending order $\hat{\pi}_{1}^{(-S(i))}(X_{i})\geq\ldots\geq\hat{\pi}_{C}^{(-S(i))}(X_{i})$, the APS score is $E_{i}=E(X_{i},Y_{i},\hat{\pi}^{-S(i)})$, where the generalized inverse quartile $E$ is defined in (4). This process yields a set of $n$ conformity scores $\{E_{1},\ldots,E_{n}\}$.
    4. (d)
      Prediction Set Construction. For a new test point $X_{n+1}$ and a desired miscoverage rate $\alpha$, we construct the prediction set $\mathcal{C}(X_{n+1})$ as follows: With the conformity scores $E_{1},\ldots,E_{n}$ obtained from the previous step, the $p$-value for a candidate class $y$ is given by:
      $\hat{p}(y)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left\{E_{i}\geq E(X_{n+1},y,\hat{\pi}^{(-S(i))})\right\}.$ (5)
      The prediction set is then formed by including all classes whose $p$-value is greater than or equal to $\alpha$:
      $\mathcal{C}(X_{n+1})=\{y\in\{1,\ldots,C\}\mid\hat{p}(y)\geq\alpha\}.$ (6)
  3. 3.
    Jackknife+-after-Bootstrap (J+ab) with APS Score [6]. The J+ab method is similar to CV+ but utilizes bootstrapping instead of $K$-fold cross-validation to create leave-one-out models and scores. This approach is particularly well-suited for ensemble methods like random forests.
    1. (a)
      Bootstrap Ensemble. We generate $B$ bootstrap samples from the training data $\{(X_{i},Y_{i})\}_{i=1}^{n}$. For each bootstrap sample $b\in\{1,\ldots,B\}$, train a classification model $\hat{\pi}^{(b)}$.
      As with the J+ab for regression, to ensure a valid finite-sample coverage guarantee, the number of bootstrap samples, $B$, must be randomized according to (3).
    2. (b)
      Leave-one-out Predictions and Scores. For each training point $(X_{i},Y_{i})$, we create an aggregated leave-one-out predictor, $\hat{\pi}^{(-i)}$, by averaging the predictions of all models $\hat{\pi}^{(b)}$ that were trained on bootstrap samples that did not include the point $(X_{i},Y_{i})$. Using this leave-one-out predictor, calculate the APS score $E_{i}=E(X_{i},Y_{i},\hat{\pi}^{(-i)})$, where the generalized inverse quartile $E$ is defined in (4).
    3. (c)
      Prediction Set Construction. For a new test point $X_{n+1}$, the prediction set is constructed by calculating a $p$-value using (5), but with $\hat{\pi}^{(-i)}$ replacing $\hat{\pi}^{(-S(i))}$:
      $\hat{p}(y)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left\{E_{i}\geq E(X_{n+1},y,\hat{\pi}^{(-i)})\right\}.$
      The final prediction set is then formed by including all classes with a sufficiently high $p$-value, as in (6):
      $\mathcal{C}(X_{n+1})=\{y\in\{1,\ldots,C\}\mid\hat{p}(y)\geq\alpha\}.$
Regularized Adaptive Prediction Sets (RAPS) [14]. This score was proposed for classification with a lot of classes (say 1000+ classes) where the ordering of the predicted probabilities does not reflect the true order. To this end, the authors propose to add to the APS’s score function a regularization term that penalizes classes with small predicted probabilities (assuming $\hat{\pi}_{1}(x)\geq\hat{\pi}_{2}(x)\geq\ldots\geq\hat{\pi}_{C}(x)$ as before):
@@
As with APS, we can use this score function for split conformal, CV+, or J+ab.
@@
Split + (R)APS $1-\alpha$ $\approx 1-\alpha$ $O(Tdn\log n)$ $O(Tn_{\operatorname{test}}\log n)$
CV+ + (R)APS $1-2\alpha-\epsilon_{K,n}$ $\gtrsim 1-\alpha$ $O(KTdn\log n)$ $O(Cn*n_{\operatorname{test}})$
J+ab + (R)APS $1-2\alpha$ $\gtrsim 1-\alpha$ $O(Tdn\log n)$ $O(Cn*n_{\operatorname{test}})$
Table 1: Comparison of coverage guarantees and computational costs for different conformal random forest methods. Here, the notation $\gtrsim 1-\alpha$ means “sometimes worse than $1-\alpha$”, $n$ is the number of training samples, $n_{\operatorname{test}}$ is the number of test samples, $T$ is the number of trees, $C$ is the number of classes, $K$ is the number of folds in CV+, and $\epsilon_{K,n}=\min\left\{2(1-1/K)/(n/K+1),(1-K/n)/(K+1)\right\}$. We assume that $T\ll n$ and $C\leq T$.
In Table 1, we compare coverage guarantees and computational costs across different conformal prediction methods for random forests. While CV+ and J+ab theoretically guarantee coverages of $\approx 1-2\alpha$, they typically achieve $\geq 1-\alpha$ coverage in practice. One downside of CV+ and J+ab with the APS and RAPS scores is the notably long prediction times due to the pairwise comparisons between train and test conformity scores across all classes, which contribute to $Cn*n_{\text{test}}$ computational cost.
@@ 2.1 Related work
The coverforest package implements conformal predictions [3, 4, 15, 1, 2] using random forest as the underlying model [7]. For similar work, conformity scores of bootstrapped regression based on iterative reweighted least squares is briefly discussed in [4]. For random forest classification, two conformity scores based on out-of-bag predictions were proposed in [16]. However, a finite-sample coverage guarantee cannot be obtained as they do not preserve exchangability [6].
@@
For classification, given a model that outputs probability estimates $\hat{\pi}_{j}(x)\approx P(Y=j|X=x)$, a basic conformal prediction method can be devised based on a score function $s(x,y)=1-\hat{\pi}_{y}(x)$. The prediction set for a new point $X_{n+1}$ is then given by $\mathcal{C}(X_{n+1})=\left\{y:s(X_{n+1},y)\leq q_{\alpha}\right\}$, where $q_{\alpha}$ is the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of the calibration scores [24, 4]. However, this method tend to outputs an empty set when the model is uncertain about the prediction, for example, $q_{\alpha}=0.2$ and the predicted probabilities of two classes are 0.5 each. This issue can be remedied by instead using the cumulative probability as in the APS method [5]. Our future plans include implementing multilabel classification with finite-sample coverage guarantee via conformal risk control [25, 26, 27].
There are several software packages that implement various types of conformal predictions, such as inductive conformal prediction and mondrian conformal prediction. However, coverforest is the first package to implement CV+, J+ab and split conformal with both APS and RAPS scores for set predictions tasks. See Table 1 for a feature comparison between coverforest and other conformal prediction packages. We will also see in Section 5 that coverforest’s implementations of APS and RAPS-based set predictions are faster than the alternative (MAPIE). by 2 to 5 times
Score Method
Package APS RAPS CV+ J+ab Split Inductive Mondrian
nonconformist [28]
Orange3-Conformal [29]
conforest [17]
crepes [30]
MAPIE [31]
coverforest
Table 2: Feature comparison between various conformal prediction software packages for set prediction tasks.
@@ 3.1 API description
For moderately large training and test sets, calculating the conformity scores across all pairs of training and test samples can be quite expensive. We contribute to performance improvement by implementing cross-calculations in Cython with parallel processing enabled. Specifically, the _giqs.pyx script in the package contains optimized code that speeds up calculation of cross-conformity scores, resulting in substantially faster prediction time than the previous package. Users can specify the number of threads through the num_threads parameter in the predict() method.
@@
3.2 Computational complexity and memory usage
As summarized in Table 1, both CV+ and J+ab methods require $Cn*n_{\text{test}}$ memory and computations for the calculation of conformity scores. To reduce the memory, the user could split the test set into several chunks and make predictions on each chunk. Alternatively, the user could use the split conformal method which requires only $n*n_{\text{classes}}$ memory and $Tn_{\text{test}}\log n$ computations while being able to achieve satisfactorily small set sizes on a large dataset.